Write equation in standard form.
The line through (-2,4)that is perpendicularto the x-axis

Answer by InYourEyes(5)

First of all you need to find the slope. Since the line is perpendicular to the x-axis there is no slope. Using the formula y-y1=m(x-x1) we can plug in the numbers.
1st method:
y-y1=m(x-x1)
y-4=m(x--2)
y-4=m(x+2)
y-4=0(x+2)
y-4=0
y=4

Or 2nd method:
standard form: y=mx+b
y=0(-2)+4
y=4
Hope i helped :)